# Cover Learning for Large-Scale Topology Representation

### v1

@@ Appendix C The filtered simplicial complex associated to a fuzzy cover
In order to put this filtered complex in the same footing as geometric complexes, which are usually indexed by the non-negative real numbers $[0,\infty)$, we apply the following standard rescaling. Consider the isomorphism of posets
$\big{(}\,\,[0,\infty)\,,\,\leq\big{)}$ $\to\big{(}\,\,(0,1]\,,\,\geq\big{)}$
@@
Appendix D Unsuitability of 1D Mapper for higher-dimensional topological inference
@@
What this means is that the nerve of the 1D Mapper cover, by design, only contains 00- and $1$-dimensional information, that is, it can detect cluster structure and loops in data, but it cannot be used to, e.g., estimate higher-dimensional homology.
Appendix E Witness complex and Ball Mapper
@@
Thus, the filtration output by the multiscale Ball Mapper algorithm of (Dłotko, 2019) consists of a finite set of filtration values of the Witness complex with parameter $v=0$.
Appendix F More on loss estimation
F.1 Loss normalization
@@
F.2 Regularization bounds geometry loss
@@
Appendix G More on implementation and examples
G.1 Software
G.1.1 Implementation
@@ G.2 Datasets
This is the dataset of (Packer et al., 2019), which we preprocess using Monocle 3 as in (Monocle, 3). This results in a point cloud with $6188$ points in $50$ dimensions. There are $28$ cell types. To help visualization, we keep $16$ cell types (those whose flare structure is clearly identified by UMAP’s 2D projection; see Figure 16), which we label from $1$ to $16$, and label the rest of the cell types by 00; see Table 3.
@@
G.3 Details on efficient topological representation experiment
@@
  • (Witness) Witness complex (De Silva & Carlsson, 2004). Recall from Appendix E that, when $v=0$, the Witness complex is essentially equivalent to multiscale Ball Mapper (Dłotko, 2019) (MS BMapper).
@@
G.4 Homology recovery quotient
@@
|$\frac{|\{r\in\mathbb{R}:\dim\mathsf{H}_{i}K_{r}=\beta_{i},\forall 0\leq i\leq n\}|}{b-a},$
@@
G.5 Why not use Mapper in the topological inference experiment
As explained in Sections 2 and D, 1D Mapper does not contain higher-dimensional information. Higher-dimensional extensions of Mapper exist, but the parameters are difficult to tune, and require higher-dimensional filter functions. Mapper is also static, that is, it produces a single simplicial complex. In order to produce a sequence of simplicial complexes, and thus a barcode that can be compared with standard approaches to topological inference, one would need to use multiscale Mapper. But to the best of our knowledge there is no implementation of a multiscale, higher-dimensional Mapper. We expect the parameters of such a procedure to be very hard to tune; this could be alleviated by using the framework of (Oulhaj et al., 2024), to produce a differentiable, multiscale, higher-dimensional Mapper.
G.6 Details on toroidal topology in neuroscience experiment
@@
[…]
@@
We also run furthest point subsaple + Vietoris–Rips on the dataset. In our experience, it is not possible to uncover the toroidal topology by directly using Vietoris–Rips: we subsampled $5000$ points and computed the Vietoris–Rips barcode, which resulted in a homology recovery quotient of 00, that is, in a filtration which at no point has toroidal topology. This agrees with the methodology of (Gardner et al., 2022), where the toroidal topology was uncovered with persistent homology only after a non-linear dimensionality reduction step using UMAP.
@@
G.7 Plotting the nerve of a cover
@@
G.8 Differentiable Mapper on visualization experiments
@@
[…]
@@
Figure 11: Ball Mapper on human data with parameter (left-to-right): $\varepsilon=0.25$, $\varepsilon=0.2$, $\varepsilon=0.15$. The cover sizes are $18$, $26$, $48$, respectively. We see that the $\varepsilon$ parameter depends significantly on the data, and is hard to tune. The underlying topology (that of a $2$-sphere) is also hard to recover: none of the simplicial complexes have the underlying topology, which agrees with the findings in Table 1 (row “Witness ($v=0$) / MS BMapper”).
@@
Table 4: Times (in seconds) taken to run ShapeDiscover on the datasets of the examples. Columns correspond to datasets together with the n_cov parameter; rows to ShapeDiscover’s subroutines. We see that the bulk of the time is taken by the optimization step, and that the topology loss accounts for a large proportion of this.

### v2

@@ Appendix D The filtered simplicial complex associated to a fuzzy cover
In order to put this filtered complex in the same footing as geometric complexes, which are usually indexed by the non-negative real numbers $0,\infty)$, we apply the following standard rescaling. Consider the isomorphism of posets
[…]
@@
Appendix E Cover learning as a coarse-graining method
Cover learning can be interpreted as a coarse-graining method, in the sense that it provides a coarse representation of geometric data. This connects our work with methods other than those in Section 2, and here we comment on this.
In (Brugnone et al., 2019) and (Huguet et al., 2023), coarsening is done by simplifying the global structure of the point cloud; for instance, by emphasizing cluster structure. The main difference with cover learning is that cover learning produces a graph (or simplicial complex) which encodes the global structure of the point cloud, and which does not have the data points as vertices (but rather groups of data points).
(Pascucci et al., 2007) uses a technique based on the Reeb graph, which is also the main motivation for Mapper, but unlike Mapper it operates on a simplicial complex (or mesh). The main difference with general cover learning is that, since they approximate a Reeb graph, they produce a one-dimensional simplicial complex, and thus, like Mapper (and as explained in Section 2 and Appendix F) it cannot be used to do higher dimensional topological inference.
In (Wolf et al., 2019), the authors present an end-to-end method for single-cell RNA-seq; the relevant part of their method is described in their section “Graph partitioning and abstraction” (pp. 7), where they explain how they coarsen an initial knn graph to a smaller graph; this is done by computing a clustering of the original graph, and then adding edges between clusters using a measure of connectivity between different clusters. For visualization purposes, their output serves very similar purposes as ours (e.g., Figure 5). Since it is not a goal of theirs, their method is not suitable for higher dimensional topological inference.
Appendix F Unsuitability of 1D Mapper for higher-dimensional topological inference
@@
What this means is that the nerve of the 1D Mapper cover, by design, only contains 0- and $1$-dimensional information, that is, it can detect cluster structure and loops in data, but it cannot be used to, e.g., estimate higher-dimensional homology.
Appendix G Witness complex and Ball Mapper
@@
Thus, the filtration output by the multiscale Ball Mapper algorithm of (Dłotko, 2019a) consists of a finite set of filtration values of the Witness complex with parameter $v=0$.
Appendix H More on loss estimation
H.1 Loss normalization
@@
H.2 Regularization bounds geometry loss
@@
Appendix I More on implementation and examples
I.1 Software
I.1.1 Implementation
@@ I.2 Datasets
This is the dataset of (Packer et al., 2019), which we preprocess using Monocle 3 as in (Monocle, 3). This results in a point cloud with $6188$ points in $50$ dimensions. There are $28$ cell types. To help visualization, we keep $16$ cell types (those whose flare structure is clearly identified by UMAP’s 2D projection; see Figure 16), which we label from $1$ to $16$, and label the rest of the cell types by 0; see Table 3.
@@
I.3 Details on efficient topological representation experiment
@@
  • (Witness) Witness complex (De Silva & Carlsson, 2004). Recall from Appendix G that, when $v=0$, the Witness complex is essentially equivalent to multiscale Ball Mapper (Dłotko, 2019a) (MS BMapper).
@@
I.4 Homology recovery quotient
@@
$\frac{|\{r\in\mathbb{R}:\dim\mathsf{H}_{i}(K_{r})=\beta_{i},\forall 0\leq i\leq n\}|}{b-a}\,$,
@@
I.5 Why not use Mapper in the topological inference experiment
As explained in Sections 2 and F, 1D Mapper does not contain higher-dimensional information. Higher-dimensional extensions of Mapper exist, but the parameters are difficult to tune, and require higher-dimensional filter functions. Mapper is also static, that is, it produces a single simplicial complex. In order to produce a sequence of simplicial complexes, and thus a barcode that can be compared with standard approaches to topological inference, one would need to use multiscale Mapper. But to the best of our knowledge there is no implementation of a multiscale, higher-dimensional Mapper. We expect the parameters of such a procedure to be very hard to tune; this could be alleviated by using the framework of (Oulhaj et al., 2024), to produce a differentiable, multiscale, higher-dimensional Mapper.
I.6 Details on toroidal topology in neuroscience experiment
@@
We run ShapeDiscover on the data of (Gardner et al., 2022) (see Section I.2 for details), with $\texttt{n\_cov}=15$; resulting in a filtered complex with $15$ vertices. The barcode thus obtained is in Figure 6, and the homology recovery quotient is $0.2$, which is remarkable given that we only used a cover with $15$ elements.
@@
We also run furthest point subsaple + Vietoris–Rips on the dataset. In our experience, it is not possible to uncover the toroidal topology by directly using Vietoris–Rips: we subsampled $5000$ points and computed the Vietoris–Rips barcode, which resulted in a homology recovery quotient of 0, that is, in a filtration which at no point has toroidal topology. This agrees with the methodology of (Gardner et al., 2022), where the toroidal topology was uncovered with persistent homology only after a non-linear dimensionality reduction step using UMAP.
@@
I.7 Plotting the nerve of a cover
@@
I.8 Differentiable Mapper on visualization experiments
@@
I.9 Other analyses
See Table 5 for an ablation study, Figure 18 for a parameter sensitivity analysis, and Figure 19 for a noise sensitivity analysis.
@@ I.9 Other analyses
Figure 8: Cover of the octopus data obtained with ShapeDiscover with default parameters and $\texttt{n\_cov}=20$.
@@
Figure 11: Ball Mapper on human data with parameter (left-to-right): $\varepsilon=0.25$, $\varepsilon=0.2$, $\varepsilon=0.15$. The cover sizes are $18$, $26$, $48$, respectively. We see that the $\varepsilon$ parameter depends significantly on the data, and is hard to tune. The underlying topology (that of a 22-sphere) is also hard to recover: none of the simplicial complexes have the underlying topology, which agrees with the findings in Table 1 (row “Witness ($v=0$) / MS BMapper”).
@@
Table 4: Times (in seconds) taken to run ShapeDiscover on the datasets of the examples. Columns correspond to datasets together with the n_cov parameter of Table 1. Rows correspond to ShapeDiscover’s subroutines. We see that the bulk of the time is taken by the optimization step, and that the topology loss accounts for a large proportion of this.
@@
Table 5: Rows correspond to datasets together with the n_cov parameter of Table 1. Columns correspond to components of the pipeline that are turned off. Entries are the output homology recovery quotient. We observe that the measure and regularization losses are essential. The initialization using clustering is also essential, at least for real datasets. The geometry loss has no substantial effect, at least for these datasets. The topology loss has an effect, but only when the initialization is random.
[…]
autotabularcenter=results/ablation.csv,
(a) 2-sph
(b) 3-sph
(c) 2-sph
(d) 3-sph
Figure 18: We run ShapeDiscover on the 2- and 3-sphere datasets, with parameters as in Table 1, except for the ones being displayed on the axes, and compute homology recovery quotient (entries).
(a) 2-sph 5
(b) 3-sph 19
Figure 19: We run ShapeDiscover on the 2- and 3-sphere datasets, with parameters as in Table 1, and compute homology recovery quotient. Data is normalized to exactly fit in a unit hypercube. Noise is additive, meaning that each coordinate of each data point is perturbed by a uniform random number in $[0,1]$. We display mean and standard deviation over $10$ runs for each level of noise.